OpenStudy (anonymous):

Given u = (5, - 4) and v = (2, 1) find u(v)

OpenStudy (anonymous):

A. 14 B. 10 C. 6 D. -3 E. 13

jimthompson5910 (jim_thompson5910):

Are those vectors?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

and you want the dot product?

OpenStudy (anonymous):

I need to find what u times v is when u is (5,4) and v is (2,1)

jimthompson5910 (jim_thompson5910):

have you ever heard of the term dot product?

OpenStudy (anonymous):

yes, i'm not quite sure how to find it though

jimthompson5910 (jim_thompson5910):

You just multiply the corresponding components. Then you add those products.

jimthompson5910 (jim_thompson5910):

Ex: (2,3) dot (4,5) = 2*4 + 3*5 = 8 + 15 = 23

OpenStudy (anonymous):

so do i just do that for (5,-4) and (2,1) and get 5*2+-4*1= 6?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Thanks so much for your help!

jimthompson5910 (jim_thompson5910):

yw